If a and b are rational and p ,q are the roots of x^2+2ax+b=0 ,then the equation with rational coefficient one of whose roots is p+q+^√p^2+q^2 is

Grade:

1 Answers

Arun
4 years ago
Dear student
 
p + q = – 2a
pq = b
 
let Y = p+q+√p^2+q^2
 
(Y + 2a)^2 = (p + q)^2 – 2 pq = 4 p^2 – 2b
Y^2 + 4ay + 2b = 0
 
hence
x^2 + 4ax + 2b = 0